[Brahms-l] Shift report 20050130 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jan 30 2005 - 15:53:27 EST
Got the first store at about 9:00
Discovered that trig 3 was counting, so I put it into the stream.
TRMRS looks good.  See e-log for more details.
First store lasted only one hour before dumping because of loss of
permit.
Moved FS to 12deg and changed field to A1219.  Also changed polarity
of D5 to B1050
Beam restored at 13:00
List of runs achieved:
Run    tr2      tr3
FS 8 deg Full field; MRS 45 deg A1050
12687  207k    82k
FS 12 deg A1219; MRS 45 deg B1050
12692  315k    43k
12693  251k    27k
12695 in progress at shift end.
Had a 45 min run while they were doing gap cleaning (12694)
Beam projected to be dumped at 17:45

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 30 15:53:46 2005

This archive was generated by hypermail 2.1.8 : Sun Jan 30 2005 - 15:54:03 EST